Green-weakness partial color blindness, or deuteranomaly, is a sex-linked recessive trait. The green-weakness allele, g, is recessive to the allele for normal color vision, G. Another form of abnormal vision is total color blindness which is due to an autosomal gene, t, which is recessive to its counterpart allele, T, for normal vision. A man with a normal phenotype marries a woman with green-weakness partial color blindness. Both individuals are heterozygous for autosomal color blindness. What phenotypes and phenotypic frequencies would be expected among their daughters and among their sons?   
Answer #1

Answer:

Man’s genotype = XGY; Tt

Woman’s genotype = XgXg; Tt

XGY; Tt (man) x (woman) XgXg; Tt ---Parents

XgT

Xgt

XGT

XGXg; TT (normal vision daughter)

XGXg; Tt (normal vision daughter)

XGt

XGXg; Tt (normal vision daughter)

XGXg; tt (abnormal vision daughter)

YT

XgY; TT (Green-weakness partial color blindness- son)

XgY; Tt (Green-weakness partial color blindness- son)

Yt

XgY; Tt (Green-weakness partial color blindness- son)

XgY; tt (Green-weakness partial color blindness & abnormal vision- son)

Normal vision daughters = 3/8

Abnormal vision daughters = 1/8

Green-weakness partial color blindness sons = 3/8

Green-weakness partial color blindness & abnormal vision sons = 1/8